# Vacuum energy in effective field theory of general relativity

### v1

@@ Abstract
In modern view, Einstein’s theory of general relativity is regarded as a leading-order approximation to an effective field theory (EFT). It is widely accepted that at low energies, all fundamental interactions including gravity can be described by an EFT Weinberg:1995mt . Not every background of general relativity represents a valid vacuum in quantum theory. Various considerations suggest that among the three maximally-symmetric backgrounds, Minkowski is the only consistent vacuum with the possibility of non-trivial cosmological history Dvali:2013eja ; Dvali:2014gua ; Dvali:2017eba ; Dvali:2020etd ; Dvali:2024dlb . In the EFT framework, the cosmological constant is just one of the parameters of the effective Lagrangian of interacting gravitational and matter fields Donoghue:1994dn ; Donoghue:2015hwa ; Donoghue:2017pgk . We seem to be rather far from constructing/discovering the fundamental theory underlying this effective theory (if it exists at all), however, we can impose conditions on parameters of the low-energy EFT such that the vacuum energy is exactly zero. Even if such an underlying fundamental theory does not exist, the property of vanishing vacuum energy can equally well be realized order-by-order in perturbation theory within the low-energy EFT. In the realm of perturbation theory, this uniquely fixes the cosmological constant as a function of other parameters of the effective Lagrangian. On the other hand, the cosmological constant gets fixed from the consistency condition for the considered EFT in Minkowski background. In particular, the consistency of an EFT in Minkoswski background, enforced by demanding the absence of the massive ghost graviton degrees of freedom, was shown to uniquely determine the cosmological constant to all orders in the loop expansion for the case of an Abelian gauge theory with spontaneous symmetry breaking coupled to the metric field Burns:2014bva . At the two-loop level, the conditions of vanishing vacuum energy and the absence of massive ghost modes were found to yield the same expressions for the cosmological constant, thanks to non-trivial cancellations between different contributions Gegelia:2019fjx ; Gegelia:2019zrv ; Epelbaum:2024qdx . Notice that while the authors of Ref. Burns:2014bva consider a particular Abelian model for demonstration, their derivation of the Ward identity and the graviton low-energy theorem, used to obtain the condition of self-consistency, applies to any local quantum field theory which is invariant under transformations of diffeomorphism, including the Standard Model (SM) coupled to general relativity. Below we give a general argument imposing that the two-loop-order results of Refs. Gegelia:2019fjx ; Gegelia:2019zrv ; Epelbaum:2024qdx hold to all orders of perturbation theory in the EFT of general relativity. That is, the consistency of the EFT in Minkowski background and the requirement of the vanishing vacuum energy lead to the same value of the cosmological constant as a parameter of the effective Lagrangian. Notice that as all $\theta$-vacuum states are degenerate in energy Jackiw:1976pf we do not expect that non-trivial structure of the vacuum affects our arguments. A more detailed treatment of this non-perturbative issue is beyond the scope of the current paper.
@@
Using the expansion around the Minkowski background and employing dimensional regularization, the first term in Eq. (13) involving $\delta(0)$ vanishes. The second term with derivatives also vanishes due to the translational invariance of the theory in a Minkowski background. The remaining third term corresponds to the vacuum energy, which hence turns out to vanish in all orders of perturbation theory.
@@ Acknowledgements.
This work was supported in part by the MKW NRW under the funding code NW21-024-A, by the Georgian Shota Rustaveli National Science Foundation (Grant No. FR-23-856), by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 885150), by CAS through a President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022), and by the EU Horizon 2020 research and innovation programme (STRONG-2020, grant agreement No. 824093).

### v2

@@ Abstract
In modern view, Einstein’s theory of general relativity is regarded as a leading-order approximation to an effective field theory (EFT). It is widely accepted that at low energies, all fundamental interactions including gravity can be described by an EFT Weinberg:1995mt . Not every background of general relativity represents a valid vacuum in quantum theory. Various considerations suggest that among the three maximally-symmetric backgrounds, Minkowski is the only consistent vacuum with the possibility of non-trivial cosmological history Dvali:2013eja ; Dvali:2014gua ; Dvali:2017eba ; Dvali:2020etd ; Dvali:2024dlb . In the EFT framework, the cosmological constant is just one of the parameters of the effective Lagrangian of interacting gravitational and matter fields Donoghue:1994dn ; Donoghue:2015hwa ; Donoghue:2017pgk . We seem to be rather far from constructing/discovering the fundamental theory underlying this effective theory (if it exists at all), however, we can impose conditions on parameters of the low-energy EFT such that the vacuum energy is exactly zero. Even if such an underlying fundamental theory does not exist, the property of vanishing vacuum energy can equally well be realized order-by-order in perturbation theory within the low-energy EFT. In the realm of perturbation theory, this uniquely fixes the cosmological constant as a function of other parameters of the effective Lagrangian. On the other hand, the cosmological constant gets fixed from the consistency condition for the considered EFT in Minkowski background. In particular, the consistency of an EFT in Minkoswski background, enforced by demanding the absence of the massive ghost graviton degrees of freedom, was shown to uniquely determine the cosmological constant to all orders in the loop expansion for the case of an Abelian gauge theory with spontaneous symmetry breaking coupled to the metric field Burns:2014bva . At the two-loop level, the conditions of vanishing vacuum energy and the absence of massive ghost modes were found to yield the same expressions for the cosmological constant, thanks to non-trivial cancellations between different contributions Gegelia:2019fjx ; Gegelia:2019zrv ; Epelbaum:2024qdx . Notice that while the authors of Ref. Burns:2014bva consider a particular Abelian model for demonstration, their derivation of the Ward identity and the graviton low-energy theorem, used to obtain the condition of self-consistency, applies to any local quantum field theory which is invariant under transformations of diffeomorphism, including the Standard Model (SM) coupled to general relativity. Below we give a general argument imposing that the two-loop-order results of Refs. Gegelia:2019fjx ; Gegelia:2019zrv ; Epelbaum:2024qdx hold to all orders of perturbation theory in the EFT of general relativity. That is, the consistency of the EFT in Minkowski background and the requirement of the vanishing vacuum energy lead to the same value of the cosmological constant as a parameter of the effective Lagrangian.
@@
Using the expansion around the Minkowski background and employing dimensional regularization, the first term in Eq. (13) involving $\delta(0)$ vanishes. In the second term the derivatives can be taken out of the path integral. Calculating the integral by perturbative expansion around the Minkowski background one obtains result which does not depend on the spacetime coordinates due to translational invariance of the Minkowski background. By acting with derivatives on this constant expression one gets vanishing result. The remaining third term corresponds to the vacuum energy, which hence turns out to vanish in all orders of perturbation theory.
@@ Acknowledgements.
We thank O. Sakhelashvili for useful comments on the manuscript. This work was supported in part by the MKW NRW under the funding code NW21-024-A, by the Georgian Shota Rustaveli National Science Foundation (Grant No. FR-23-856), by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 885150), by CAS through a President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022), and by the EU Horizon 2020 research and innovation programme (STRONG-2020, grant agreement No. 824093).